### v1

# To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States

Abstract
The adoption of electric vehicles (EVs) is considered critical to achieving climate goals, yet it hinges on consumer interest. This study explores how public intent to purchase EVs relates to four unexamined factors: exposure to EV information, perceptions of EVs’ environmental benefits, views on government climate policy, and confidence in future EV infrastructure; while controlling for prior EV ownership, political affiliation, and demographic characteristics (e.g., age, gender, education, and geographic location). We utilize data from three nationally representative opinion polls conducted by the Pew Research Center between 2021 and 2023, and employ Bayesian techniques to estimate the ordinal probit and ordinal quantile models. Results from ordinal probit show that respondents who are well-informed about EVs, perceive them as environmentally beneficial, or are confident in development of charging stations are more likely to express strong interest in buying an EV, with covariate effects--a metric rarely reported in EV research--of 10.2, 15.5, and 19.1 percentage points, respectively. In contrast, those skeptical of government climate initiatives are more likely to express no interest, by more than 10 percentage points. Prior EV ownership exhibits the highest covariate effect (ranging from 19.0 to 23.1 percentage points), and the impact of most demographic variables is consistent with existing studies. The ordinal quantile models demonstrate significant variation in covariate effects across the distribution of EV purchase intent, offering insights beyond the ordinal probit model. This article is the first to use quantile modeling to reveal how covariate effects differ significantly throughout the spectrum of EV purchase intent.
keywords:
Decarbonization, electric vehicle, ordinal probit, Pew Research, quantile regression, technology adoption.
journal: arXiv.org
1 Introduction
Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half (Environmental Protection Agency, 2024). Electric vehicles (EVs) are widely regarded as a key solution for reducing emissions and meeting climate targets, given their zero tailpipe emissions. To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives. In the US, key measures include nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program..
Despite these efforts, EV sales in the US remain relatively low, accounting for only about 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024). At the same time, the future trajectory of EV adoption in the US is increasingly uncertain. In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand (Colias and Otts, 2024). Moreover, successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations. The Biden administration has promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its climate agenda, while the Trump administration has signaled a deregulatory shift on EV policies with plans to weaken both incentives as well as regulations. [^2] President Trump’s Executive Order on January 20, 2025, outlined plans to halt \$7.5 billion in charging infrastructure funding, weaken vehicle emissions standards that drive EV production, eliminate EV tax credits, and revoke California’s EV mandate waiver, which would limit state-level regulations. See https://www.whitehouse.gov/presidential-actions/2025/01/unleashing-american-energy/ These shifting political priorities may reflect broader tensions between environmental regulations, industry interests, and fiscal considerations in EV adoption. While EVs help reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024). The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and restrict consumer choice.
Ultimately, the future of the EV market depends on consumer interest, which is likely to be shaped by multiple factors, including cost considerations, convenience, perception of environmental benefits, and awareness of these trade-offs. Although EVs often have higher upfront costs due to expensive battery technology, they can offer long-term savings through lower operational and maintenance expenses, as electric motors require less maintenance. However, the overall cost of owning an EV can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024). Beyond cost considerations, access to reliable charging infrastructure remains an important consideration. Limited charging availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel. Environmental concerns also shape consumer attitudes, but perceptions may vary, given the regional heterogeneity in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
Recognizing the importance of public interest in shaping the future of this market, this study examines how public opinion on purchasing EVs is associated with factors such as exposure to EV information, perceptions about the environmental impact of EVs (as compared to gas-powered vehicles) and government actions to mitigate climate change, confidence in EV related infrastructure development, prior ownership, and demographic characteristics such as age, gender, education, political affiliation, and geographical location of the respondents.
Our paper makes three key contributions. First, we explore the role of several novel factors that have been largely overlooked in the existing EV adoption literature. Previous studies have examined how a combination of factors, including range anxiety, performance concerns, charging infrastructure, environmental attitudes, government policies, peer effects, and demographic characteristics, influences EV purchase intent (Carley et al., 2013; Tiwari et al., 2020; Zhao et al., 2022; Mamkhezri, 2025). Other studies examine the impact of specific factors, such as providing information about total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), reputation (Buhmann and Criado, 2023), and consumer incentives (Xue et al., 2023; Stekelberg and Vance, 2024). Related research examines factors driving actual EV sales rather than purchase intent, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023). We contribute to this literature by examining how EV purchase intent is influenced by factors such as confidence in future charging infrastructure development, perceptions of government climate efforts, exposure to EV-related information, and perceptions of EVs’ environmental benefits.
Our second contribution is related to usage of large and nationally representative datasets. Existing research studying the trends and determinants of public opinion and purchase intent for EVs has predominantly relied on small-scale surveys. For example, Tiwari et al. (2020) analyzed 1,800 responses from UK residents, Lashari et al. (2021) analyzed 1,500 responses from South Korea, Hu et al. (2023) analyzed 807 responses from China, and Mamkhezri (2025) analyzed 1,500 responses from the US. [^3] While some recent studies (e.g., Pani et al. (2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of perception. Singh et al. (2020) provide a meta-analysis of 211 peer-reviewed research articles published between 2009 and 2019, identifying key factors influencing public opinion of EVs. Notably, most such studies are either based on a single survey for a particular year, not nationally representative, or constrained by sample size; which diminishes their generalizability. This paper overcomes such data related limitations by drawing on three nationally representative opinion polls conducted by the Pew Research Center in the US with a significantly larger sample size than existing previous studies. Specifically, we use data from 2021, 2022, and 2023, with 11,052, 7,173, and 7,201 observations, respectively.
Our third contribution lies in the application of advanced modeling and estimation techniques to study EV purchase intent. Existing research has used various methods to analyze EV purchase intent, such as linear regression to assess the effects of personal beliefs and preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation in shaping EV preferences (Buhmann and Criado, 2023); mixed logit models to explore attitudes towards various EV policy incentives, employment changes and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022). These papers rely on Classical (Frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters. This can lead to biased estimates of covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly translate to changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018).
In contrast, this paper adopts a Bayesian modeling framework that provides a posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables us to make direct probability statements about the parameters--rather than relying on the notion of “confidence” as in the Classical framework. Our empirical analysis consists of two stages. First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how different factors are associated with consumer interest in EV adoption. We present the covariate effects on the average probabilities of outcomes, an aspect which is rarely explored in the EV literature. Next, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how different factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined. Notably, this study is the first to apply ordinal quantile analysis to EV purchase intent. Since no classical estimation methods exist for ordinal quantile models, Bayesian estimation is essential for this analysis. The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
The remainder of the article is structured as follows. Section 2 introduces the ordinal probit and ordinal quantile models employed in our analysis, along with their MCMC algorithms for estimation. Section 3 outlines the data and provides a preliminary investigation. In Sections 4, we present the results from our models and discuss their implications. Finally, Section 5 offers concluding remarks.
2 The Modeling Framework
In this section, we briefly describe the ordinal probit model followed by the ordinal quantile model. For both the models, we outline the Bayesian estimation algorithm using Markov chain Monte Carlo (MCMC) methods and explain the computation of covariate effects.
2.1 The Ordinal Probit Model
Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates. These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al. (2023). In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked. However, these scores only carry ordinal meaning and lack cardinal interpretation. Thus, while the categories are ranked, the differences between them cannot be compared directly. For example, in our current study, responses to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’. Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1. Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
The ordinal regression model, following Albert and Chib (1993), can be written in terms of a continuous latent variable $z_{i}$ as follows:
$z_{i}=x^{\prime}_{i}\beta+\varepsilon_{i},\hskip 54.2025pt\forall\;i=1,\cdots,n,$ (1)
where $x_{i}$ is a $k\times 1$ covariate vector including a column of ones, $\beta$ is a $k\times 1$ parameter vector, and $n$ is the number of observations. We assume the error $\varepsilon$ is independent and identically distributed (iid) as a standard normal distribution i.e., $\varepsilon_{i}\sim N(0,1)$ for $i=1,2,\ldots,n$; which gives rise to the ordinal probit model.
In our study, the unobserved variable $z_{i}$ represents intent to purchase EV and is associated with the observed discrete response $y_{i}$, that denotes response to the question about purchasing an EV, as follows:
[…]
where $-\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$. This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is visually represented in Figure 1. As seen from the figure, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so on. We set $\gamma_{1}=0$ and $var{(\varepsilon)}=1$ to anchor the location and scale of the distribution, respectively. These restrictions are required for parameter identification and further explained in (Jeliazkov et al., 2008) and Jeliazkov and Rahman (2012).
Figure 1: A pictorial representation of the distribution of the latent variable $z$. The four probabilities P$(y_{i}=1)$, P$(y_{i}=2)$, P$(y_{i}=3)$ and P$(y_{i}=4)$ correspond to the responses: ‘not at all likely’, ‘not too likely’, ‘somewhat likely’, and ‘very likely’, regarding the intention to purchase an electric vehicle ($z_{i}$) for individual $i$ with mean $x^{\prime}_{i}\beta$.
Given the data $y$ = $(y_{1},\cdots,y_{n})^{\prime}$ and the covariates, the likelihood function for ordinal probit model can be written as follows:
$\begin{split}f(y|\beta,\gamma)&=\prod_{i=1}^{n}\prod_{j=1}^{J}\Pr(y_{i}=j|\beta,\gamma)^{I(y_{i}=j)}=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg{[}\Phi(\gamma_{j}-x^{\prime}_{i}\beta)-\Phi(\gamma_{j-1}-x^{\prime}_{i}\beta)\bigg{]}^{I(y_{i}=j)},\end{split}$ (3)
where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise. But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling. So, we apply the transformation of Albert and Chib (2001): $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $12\leq j\leq J-$ 1, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
To complete the Bayesian framework, we utilize the Bayes theorem to combine the augmented likelihood $f(y,z|\beta,\delta)$ with the prior distributions: $\beta\sim N\left(\beta_{0},B_{0}\right)$ and $\delta\sim N\left(d_{0},D_{0}\right)$. The resulting augmented posterior distribution can be written as,
$\begin{split}\pi\left(\beta,\delta,z|y\right)&\propto f\left(y|\beta,\delta,z\right)f(z|\beta)\pi(\beta)\pi(\delta)=\left\{\prod\limits_{i=1}^{n}f\left(y_{i}|\delta,z_{i}\right)\right\}f(z|\beta)\pi(\beta)\pi(\delta),\end{split}$ (4)
where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, the correspondence between $\gamma$ and $\delta$ is determined by the one-to-one mapping, and the cutpoint index $j$ is determined by the observed value of $y_{i}$. Additionally, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$ and the priors $\pi(\beta)$ and $\pi(\delta)$ are pdf’s of normal distribution. Based on the augmented joint posterior (4), the conditional posteriors can be derived and the model can be estimated using MCMC sampling, as outlined in Algorithm 1.
Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effect of a variable. Let’s assume that the $l$-th covariate, ${x_{i,l}}$, is set to two distinct values, $a$ and $b$, denoted as ${x_{i,l}^{a}}$ and ${x_{i,l}^{b}}$, respectively. We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{p,l},\beta_{p,-l})$, where $-l$ denotes all covariates (and parameters) except for the $l$-th covariate (and parameter). Our goal is to examine the distribution of the difference ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$ for $1\leq j\leq J$, by marginalizing over ${x_{i,-l}}$ and $(\beta_{p},\delta_{p})$. The marginalization enables us to incorporate the uncertainty associated with data and model parameters, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
Algorithm 1: MCMC algorithm for estimating ordinal probit model.
  1. 1.
    Sample $\delta,z|y,\beta$ in one block as follows:
    1. (a)
      Sample $\delta|y,\beta$ marginally of $z$ by generating $\delta^{\prime}$ from a random-walk chain $\delta^{\prime}=\delta+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter, $\hat{D}=\left.-\left[\frac{\partial^{2}\ln f(y|\beta,\delta)}{\partial\delta\partial\delta^{T}}\right]^{-1}\right|_{\delta=\hat{\delta}}$, and $\hat{\delta}=\arg\max_{\delta}\ln f(y|\beta,\delta)$. Accept the proposed value $\delta^{\prime}$ with probability
      $\alpha_{MH}(\delta,\delta^{\prime})=\min\left\{1,\frac{f(y|\beta,\delta^{\prime})\pi(\delta^{\prime})}{f(y|\beta,\delta)\pi(\delta)}\right\},$
      otherwise repeat the current value $\delta$.
    2. (b)
      Sample $z_{i}|y,\beta,\delta\sim TN_{\left(\gamma_{j-1},\gamma_{j}\right)}\left(x_{i}^{\prime}\beta,1\right)$ for $i=1,\ldots,n$, where $TN_{(\gamma_{j-1},\gamma_{j})}$ denotes a truncated normal distribution constrained between $\gamma_{j-1}$ and $\gamma_{j}$; and $\gamma$ is derived from the one-to-one mapping relating $\gamma$ and $\delta$.
  2. 2.
    Sample $\beta|z\sim N(\hat{\beta},\hat{B})$ where $\hat{B}^{-1}=({B}_{0}^{-1}+X^{\prime}X)$, and $\hat{\beta}=\hat{B}(B_{0}^{-1}\beta_{0}+X^{\prime}z)$.
To obtain draws from the distribution ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$, we use the method of composition (Chib and Jeliazkov, 2006). This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference ${\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)}$, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$. The process is repeated for all individuals and MCMC draws. The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences:
$ACE_{y=j}\approx\frac{1}{M}\frac{1}{n}\sum_{m=1}^{M}\sum_{i=1}^{n}\Big{[}\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta^{(m)},\delta^{(m)})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta^{(m)},\delta^{(m)})\Big{]},$ (5)
where $(\beta^{(m)},\delta^{(m)})$ denotes the MCMC draws, and $M$ is the number of post burn-in MCMC draws.
2.2 The Ordinal Quantile Model
The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates. However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions. For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs. However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers. To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression. These insights provide suggestive evidence for more targeted interventions; for example, they show that improving infrastructure visibility may be more effective when directed at consumers already considering EVs, rather than those with little or no interest.
The ordinal quantile regression model, proposed by Rahman (2016), can be written in terms of a latent variable $z_{i}$ as follows:
$z_{i}$ $=x^{\prime}_{i}\beta_{p}+\varepsilon_{i},$ $\qquad\forall\;i=1,\cdots,n,$ (6)
[…]
where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution (Yu and Zhang, 2005). The AL distribution is essential to create a working likelihood function, and yields the following expression,
$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}P(y_{i}=j|\beta_{p},\gamma_{p})^{I(y_{i}=j)}\\ &=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg{[}F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg{]}^{I(y_{i}=j)},\end{split}$ (7)
where $F_{AL}(\cdot)$ denotes the cdf of an AL distribution and $I(y_{i}=j)$ is an indicator function.
At this stage, three key issues need attention. First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $12\leq j\leq J-$ 1, where note that $\delta_{p}$ is quantile dependent. Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is fixed for any quantile to meet the scale restriction. Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors. So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$. The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$. Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
By Bayes theorem, we combine the augmented likelihood $f(y,z|\beta_{p},\delta_{p},w)$ with the normal priors: $\beta_{p}\sim N(\beta_{p0},B_{p0})$, $\delta_{p}\sim N(\delta_{p0},D_{p0})$, to arrive at the augmented joint posterior distribution,
$\begin{split}\pi(z,\beta_{p},\delta_{p},w|y)&\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big{\{}\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big{\}}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg{\{}\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg{\}}\\ &\qquad\times\;N(\beta_{p}|\beta_{p0},B_{p0})\;N(\delta_{p}|\delta_{p0},D_{p0}).\end{split}$ (8)
In the first line, we use the factorization $f(y,z|\beta_{p},\delta_{p},w)=f(y|z,\beta_{p},\delta_{p},w)\times f(z|\beta_{p},w)$ and note that $z$ is independent of $\delta_{p}$. In the second line, we exploit the independence of $y_{i}$ from $(\beta_{p},w)$ given $(z_{i},\delta_{p})$, which follows from the second line of equation (6), where $y_{i}$ given $(z_{i},\delta_{p})$ is determined with probability 1. In the third and fourth lines, we specify the conditional distribution of the latent variable $z$ and the prior distribution on the parameters $(\beta_{p},\delta_{p})$.
The conditional posterior distributions are derived from the augmented joint posterior distribution (i.e. equation 8), and the parameters are sampled as per Algorithm 2. The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
Algorithm 2: MCMC algorithm for estimating ordinal quantile model.
  1. 1.
    Sample $\beta_{p}|z,w$ $\sim$ $N(\tilde{\beta}_{p},\tilde{B}_{p})$, where,
    $\tilde{B}^{-1}_{p}=\bigg{(}\sum_{i=1}^{n}\frac{x_{i}x^{\prime}_{i}}{\tau^{2}w_{i}}+B_{p0}^{-1}\bigg{)},\quad\mbox{and}\quad\tilde{\beta}_{p}=\tilde{B}_{p}\bigg{(}\sum_{i=1}^{n}\frac{x_{i}(z_{i}-\theta w_{i})}{\tau^{2}w_{i}}+B_{p0}^{-1}\beta_{p0}\bigg{)}.$
  2. 2.
    Sample $w_{i}|\beta_{p},z_{i}$ $\sim$ $GIG\,(0.5,\tilde{\lambda}_{i},\tilde{\eta})$, for $i=1,\cdots,n$, where,
    $\tilde{\lambda}_{i}=\Big{(}\frac{z_{i}-x^{\prime}_{i}\beta_{p}}{\tau}\Big{)}^{2},\quad\mbox{and}\quad\tilde{\eta}=\Big{(}\frac{\theta^{2}}{\tau^{2}}+2\Big{)}.$
  3. 3.
    Sample $\delta_{p}|y,\beta_{p}$ marginally of $w$ (latent weight) and $z$ (latent data), by generating $\delta_{p}^{\prime}$ from a random-walk chain $\delta^{\prime}_{p}=\delta_{p}+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter and $\hat{D}$ denotes negative inverse Hessian, obtained by maximizing the log-likelihood with respect to $\delta_{p}$. Given the current value of $\delta_{p}$, accept the proposed draw $\delta^{\prime}_{p}$ with probability,
    $\alpha_{MH}(\delta_{p},\delta^{\prime}_{p})=\min\bigg{\{}1,\frac{\;f(y|\beta_{p},\delta^{\prime}_{p})\;\pi(\beta_{p},\delta^{\prime}_{p})}{f(y|\beta_{p},\delta_{p})\;\pi(\beta_{p},\delta_{p})}\bigg{\}};$
    otherwise repeat the value $\delta_{p}$. The variance of $s$ may be tuned for appropriate step size and acceptance rate.
  4. 4.
    Sample $z_{i}|y,\beta_{p},\gamma_{p},w$ $\sim$ $TN_{(\gamma_{p,j-1},\gamma_{p,j})}(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\cdots,n$, where $TN$ denotes a truncated normal distribution and $\gamma_{p}$ is obtained via $\delta_{p}$ using the one-to-one mapping.
After estimating the ordinal quantile models, we compute the covariate effects using the method of composition. Existing studies on quantile regression that have applied this method to estimate the covariate effects include (Rahman and Vossmeyer, 2019) and Bresson et al. (2021). Assuming a setup similar to Section 2.1, our goal is to obtain draws from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$. We randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$. This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.
3 Data
The study draws on survey data from three opinion polls carried out by the Pew Research Center over a three-year span. The first survey was conducted during April 20-21, 2021 on a sample of 13,749 adults, the second survey took place during May 2-8, 2022, on a sample of 10,282 adults, and the third survey was administered during May 30-June 4, 2023 on a sample of 10,329 adults. After removing all missing observations and cleaning the data, we have a sample of 11,052, 7,173, and 7,201 observations available for analysis. Table 1 shows a descriptive summary of the independent and dependent variables. Specifically, it presents the mean and standard deviation of the continuous variables, and count and percentage for the categorical variables.
The above survey data offer at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent. First, our survey data is nationally representative and span multiple years. As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity. All survey data and additional information on the surveys used in this paper are available for download from the Pew Research Center dataset portal (and the corresponding author’s webpage). Second, an appealing feature of our data sets is the availability of ordinal responses to likelihood of purchasing EV. This enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023). Third, we note that the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase. Therefore, if there was an increase in interest for sustainable products like electric vehicles during the pandemic, this should be evident in the data summary, with a higher percentage of respondents expressing a preference for buying electric vehicles during the pandemic phase.
Table 1: Descriptive summary of the variables for three Pew Research Center surveys: April 2021; May, 2022; and May-Jun, 2023. Mean (mean) and standard deviations (std) are reported for variables treated as continuous, while count (count) and percentages (per) are reported for categorical variables.
April 2021 May 2022 May–June 2023
variable mean std mean std mean std
Income/10000 (USD) 7.00 3.33 8.00 3.23 8.00 3.24
count per count per count per
EV Info $\circ$ Read or heard a lot 4170 37.73 ..absent... . ..absent... . ..absent... . ..absent... .
$\circ$ Little or nothing at all 6882 62.27 ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $\circ$ EVs are environmentally ‘better’ than gas-powered vehicles 7944 71.88 ..absent... . ..absent... . ..absent... . ..absent... .
$\circ$ About the same or worse 3108 28.12 ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra $\circ$ US will build charging stations and infrastructure: extremely or very confident ..absent... . ..absent... . ..absent... . ..absent... . 1263 17.54
$\circ$ Somewhat confident, not too confident, or not at all confident ..absent... . ..absent... . ..absent... . ..absent... . 5938 82.46
EV Owner Owns electric or hybrid vehicle 1039 9.40 769 10.72 812 11.28
GACC $\circ$ Federal government is doing ‘too much’ to reduce the effects of climate change 1962 17.75 1684 23.48 1736 24.11
$\circ$ About the right amount 2203 19.93 1636 22.81 1582 21.97
$\circ$ Too little 6887 62.31 3853 53.72 3883 53.92
Age $<50$ Indicator for age below 50 5124 46.36 2687 37.46 2596 36.05
Female Indicator for female gender 5848 52.91 3819 53.24 3815 52.98
$\circ$ Graduate and above 6447 58.33 3846 53.62 3730 51.80
Education $\circ$ Some college 3194 28.90 2201 30.68 2276 31.61
$\circ$ HS and below 1411 12.77 1126 15.70 1195 16.59
Married Married or living with a partner ..absent... . ..absent... . 5692 79.35 5640 78.32
Metropolitan Lives in a metropolitan area 9817 88.83 6260 87.87 6286 87.29
$\circ$ Northeast 1784 16.14 1100 15.34 1084 15.05
Region $\circ$ Midwest 2444 22.11 1613 22.49 1586 22.02
$\circ$ South 4245 38.41 2841 39.61 2987 41.48
$\circ$ West 2579 23.34 1619 22.57 1544 21.44
US Born Indicator for born in the US 9664 87.44 6150 85.74 6188 85.93
$\circ$ White 7801 70.58 5335 74.38 5083 70.59
Race $\circ$ Black 820 7.42 446 6.22 726 10.08
$\circ$ Other races 2431 22.00 1392 19.41 1392 19.33
$\circ$ Democrat 4201 38.01 2258 31.48 2268 31.50
Party $\circ$ Republican 2667 24.13 2453 34.20 2341 32.51
$\circ$ Independent & others 4184 37.86 2462 34.32 2592 36.00
Very likely 2411 21.82 1526 21.27 1397 19.40
Purchase EV Somewhat likely 2819 25.51 1936 26.99 1815 25.20
Not too likely 3381 30.59 1787 24.91 1736 24.11
Not at all likely 2441 22.09 1924 26.82 2253 31.29
We begin by discussing the independent variables that have a direct influence on purchasing an EV–the question of primary interest in this article. Two such variables, ‘EV Info’ and ‘Env Better’, are specific to the April 2021 survey. The variable ‘EV Info’ measures how much a respondent has read or heard about EVs. It is coded as 1 if the respondent has read or heard a lot about EVs, and 0 if they have read or heard little or nothing. As shown in Table 1, about 38% of respondents have read or heard a lot about EVs. The variable ‘Env Better’ reflects perception about the environmental impact of EVs as compared to gas-powered vehicles. This indicator variable takes the value 1 if the respondent believes that EVs are better for the environment than gas-powered vehicles, and 0 otherwise (worse for the environment or about the same). According to Table 1, approximately 72% of respondents view EVs as being environmentally superior to gas-powered vehicles. Another noteworthy variable is ‘EV Infra’, which is only available in the May-June 2023 survey. This indicator variable is coded as 1 if the respondent is extremely or very confident that the US will build the necessary charging stations and infrastructure to support large number of EVs on the roads, and 0 otherwise (somewhat confident, not too confident, or not at all confident). As seen in Table 1, about 82% of respondents express a negative perception towards the government for creating the infrastructure needed to support the adoption of EVs.
Two key variables of interest to our study, which are available in all the surveys, are ‘EV Owner’ and ‘GACC’. The ‘EV Owner’ variable indicates whether the respondent currently owns an electric or hybrid vehicle and takes the value 1 if true and 0 otherwise. As shown in Table 1, the percentages of respondents who are current owners of an electric or hybrid vehicle are 9.40%, 10.72%, and 11.28% across the three survey datasets. The ‘GACC’ variable reflects respondents’ perceptions of government actions (GA) to mitigate global climate change (CC). This variable is divided into three categories: ‘too much’, ‘about the right amount’, and ‘too little’ (with ‘too little’ serving as the reference category). Table 1 indicates that in each sample, more than half of the respondents feel that government efforts to combat global climate change are insufficient.
We also incorporate a broad set of demographic variables as covariates in our regression models. These variables (except marital status) are common across all surveys and include: income, age, gender, education, marital status (not available in the April 2021 survey), metropolitan area, region of residence, birth country, racial background, and party affiliation. We now examine the construction of these variables in more detail.
Pew surveys record income as belonging to one of nine categories: $0-30k,30k-40k,40k-50k,50k-60k,60k-70k,70k-80k,80k-90k,90k-100k$, and $>100k$, where $k$ denotes a thousand dollars. To simplify, we use the mid-point of each income range and assign an imputed value of \$20,000 for the first category and \$110,000 for the last category. Age is recorded into four categories (18-29, 30-49, 50-64, and $\geq 65$ years), but for convenience we recode it to a binary indicator taking the value 1 if the respondent is below 50, and 0 otherwise. As shown in Table 1, 46.36% of respondents in the first dataset belong to the younger age group, with this percentage being approximately 10% lower in the following two datasets. Gender is represented by a binary indicator ‘Female’, where 1 indicates female and 0 indicates male. We see that more than 50% respondents across all three datasets are female. Education is classified into three categories: ‘Graduate & Above’, ‘Some College’, and ‘High School (HS) and Below’; the latter is used as the reference category. As shown in Table 1, the majority of respondents are graduates or have higher education, followed by those with some college education. Therefore, respondents across the three samples are generally well-educated.
Information on marital status is available only in the May 2022 and May--June 2023 surveys. We create an indicator variable ‘Married’ that takes value 1 if the respondent is married or living with a partner, and 0 if they are divorced, separated, widowed, or never been married. As per Table 1, more than three-quarters of respondents are married or living with a partner. Two variables reflect a respondent’s residential background. The ‘Metropolitan’ variable is an indicator that equals 1 if the respondent lives in a metropolitan area, and 0 otherwise. Across the three datasets, more than four-fifths of the respondents are located in metropolitan areas. Additionally, we include regional indicators for ‘Northeast’, ‘Midwest’, ‘South’, and ‘West’; with Northeast serving as the reference category in our regression models. As shown in Table 1, the largest proportion of respondents (about 40%) reside in the South, followed by approximately 22% in both the Midwest and West. The smallest fraction, about 15%, lives in the Northeast.
The variable ‘US Born’ indicates whether the respondent was born in the US or is an immigrant. According to Table 1, a predominant fraction (about 86%) of the sample across three datasets were born in the US. Our demographic variable list also includes race indicators for ‘White’ (reference category), ‘Black’, and ‘Other Races’. As expected, the largest proportion of respondents (over 70%) are White, followed by about 20% identifying as belonging to other races. The smallest group, about 8%, are Black. Finally, party affiliation is captured with indicator variables for ‘Democrat’, ‘Republican’ (reference category), and ‘Independent & Others’. Respondents are roughly equally distributed across these three political categories.
As noted earlier, our objective is to analyze the factors that influence individuals’ decision to purchase an EV. However, a continuous measure of support or opposition is not directly observable. The data is limited to the respondent’s answer to the question: “The next time you purchase a vehicle, how likely are you to seriously consider purchasing an electric vehicle?” Response to this question serves as the dependent variable in our models and has four categories: ‘very likely’, ‘somewhat likely’, ‘not too likely’, and ‘not at all likely’. The surveys also include a fifth category :‘I do not expect to purchase a vehicle’; we exclude responses in this category from our analysis to simplify the framework and adhere to ordinal modeling. Table 1 presents the count and percentage of each response category in the three datasets. We see from Table 1, 47.33% of the respondents in the April 2021 survey and 48.26% of the respondents in the May 2022 survey are either ‘very likely’ or ‘somewhat likely’ to purchasing an EV. In comparison, the corresponding percentage in the May-June 2023 survey is 44.60%. Recall, the first two surveys were conducted during the Covid-19 period, while the third survey took place in the post-Covid period. Based on these percentages, we do not observe a significant decline in preference for purchasing EVs in the post-pandemic phase. Overall, our data summary clearly shows that public attitude towards purchasing an EV is generally positive.
Figure 2: Stacked bars displaying the percentage of observations corresponding to the four opinion categories for a selected list of covariates. In each panel, the first stacked bar shows the percentage of observations for ‘not at all likely’ and the cumulative percentage for ‘not at all likely’ and ‘not too likely’. The second stacked bar shows the percentage of observations for ‘somewhat likely’ and the cumulative percentage for ‘somewhat likely’ and ‘very likely’.
Further insights into the association between the dependent variable and selected regressors from the three survey datasets can be gained from Figure 2. In each panel, the bars represent two ‘opinion’ categories: ‘not likely’ and ‘likely’. The first stacked bar displays the percentage of responses for ‘not at all likely’ (shown in dark shade) and the cumulative percentage for both ‘not at all likely’ and ‘not too likely’ (sum of dark and light shades). The second stacked bar illustrates the percentage of responses for ‘somewhat likely’ (in dark shade) and the cumulative percentage for ‘somewhat likely’ and ‘very likely’ (sum of dark and light shades).
The first two panels of Figure 2 highlight the influence of information on consumers’ decisions to purchase EVs. In the first panel, among respondents who have ‘read or heard little or nothing at all’ about EVs, only 46.99% are likely to consider purchasing one. In contrast, the second panel reveals a much higher likelihood of purchase—62.07%—among those who have ‘read or heard a lot’ about EVs. A more striking difference emerges when examining perceptions of the environmental impact of vehicles. As shown in the third panel, just 19.18% of respondents who view EVs as ‘about the same or worse’ than gas-powered vehicles are likely to purchase one. However, as shown in the fourth panel, the percentage jumps to 65.79% for those who believe EVs have a lower environmental impact. A similar pattern is evident, from panels five and six, when considering respondents’ views on the development of charging stations and infrastructure to support large number of EVs on the road. Among those who lack confidence that the government will build the necessary facilities, only 36.49% are likely to buy an EV. In stark contrast, a much higher 82.74% of respondents who believe the government will build the required facilities for EVs are likely to make a purchase. These relationships highlight how knowledge and environmental impact of EVs, and confidence in supporting infrastructure play pivotal roles in shaping consumer purchase decisions.
The remainder of panels in Figure 2 (i.e., panels seven to eighteen) exhibit how factors such as ownership of an EV or hybrid vehicle, age, and gender affect the decision to purchase an EV. The stacked bars from panels seven to twelve clearly show that owning an EV or hybrid vehicle (as opposed to not) increases the likelihood of purchasing an EV across all survey datasets. For instance, 84.12% of EV or hybrid owners are likely to purchase an EV as compared to only 49.42% (a difference of 34.7%) in the April 2021 survey data. For the same comparison, a difference of 36.98% ($81.28\%-44.3\%$) and 36.47% ($76.97\%-40.5\%$) is noted in the May 2022 and May-June 2023 survey data. Age also plays a role, with a majority of respondents under 50 being likely to purchase an EV across all three datasets. With respect to gender, the likelihood of purchasing an EV among females is 53.27% in the April 2021 dataset, decreases to 49.18% in May 2022, and further drops to 44.69% in the May-June 2023 dataset.
4 Empirical Results and Discussion
In this section, we report the results from the Bayesian estimation of ordinal probit and ordinal quantile models. We also report the marginal likelihood which may be used for model comparison. Morevoer, for each model, we compute the covariate effects for all or some of the significant variables and discuss our findings in the context of existing literature.
4.1 Results from Ordinal Probit Models
We begin by presenting the results from ordinal probit models estimated separately for each survey data. Specifically, Table 2 presents the posterior means and standard deviations of the parameters along with the logarithm of marginal likelihood (Chib and Jeliazkov, 2001). The estimates are obtained from 10,000 MCMC iterations (after a burn-in of 2,500 iterations) and utilizes the following prior distributions: $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\ast I)$, where $I$ denotes an identity covariance matrix. The priors are relatively diffuse to allow the data speak for itself. In ordinal models, the sign of coefficients indicates the direction of covariate effects for the first and last response categories. For example, in Table 2, the coefficient for “EV Info” is positive (0.43), which implies that it increases (decreases) the probability of last (first) category “Very likely” (“Not at all likely”) to purchase an EV. However, these coefficients do not represent the covariate effects nor can we state anything a-priori about the covariate effects on intermediate outcomes (i.e., “Not too likely” and “Somewhat likely”). Therefore, Table 3 provides the covariate effects for all variables which are statistically different from zero. Note, our measure of covariate effects account for uncertainty in parameters and observations to avoid biases (Jeliazkov and Vossmeyer, 2018).
The combined results from the 2021, 2022, and 2023 surveys reveal consistent patterns in factors associated with EV purchase intent. As shown in Table 3, the covariate effects corresponding to the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs as compared to those with limited exposure to EV-related information. These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories. Our findings align with Zhao et al. (2022) and Scherrer (2023), who demonstrate that greater exposure to information about EVs—whether through peer effects or media—positively influences attitudes toward EVs, and with Krause et al. (2013), who underscore the crucial role of accurate information in shaping consumer perceptions.
Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability. For the 2021 survey, we also observe, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent. This finding aligns with Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs  [^4] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
Table 2: Results from ordinal probit models: posterior mean (Mean) and standard deviation (Std) of the parameters, and logarithm of marginal likelihood, from April 2021; May 2022; and May–June 2023 survey data.
April 2021 May 2022 May–June 2023
Mean Std Mean Std Mean Std
Intercept $-0.28$ $0.07$ $0.36$ $0.09$ $-0.05$ $0.09$
EV Info $0.43$ $0.02$.absent... . ..absent... . ..absent... . ..absent... .
Env Better $0.79$ $0.03$.absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . $0.82$ $0.04$
EV Owner $0.79$ $0.04$ $0.86$ $0.05$ $0.82$ $0.04$
GACC (TM) $-0.79$ $0.04$ $-1.13$ $0.04$ $-1.07$ $0.04$
GACC (RA) $-0.52$ $0.03$ $-0.63$ $0.03$ $-0.49$ $0.04$
Age $<50$ $0.16$ $0.02$ $0.17$ $0.03$ $0.15$ $0.03$
Female $-0.02$ $0.02$ $-0.11$ $0.03$ $-0.12$ $0.03$
Grad & Above $0.38$ $0.04$ $0.45$ $0.04$ $0.49$ $0.04$
Some College $0.18$ $0.04$ $0.23$ $0.04$ $0.21$ $0.04$
Married ..absent... . ..absent... . $0.06$ $0.03$ $0.02$ $0.04$
Metropolitan $0.24$ $0.03$ $0.25$ $0.04$ $0.20$ $0.04$
Midwest $-0.08$ $0.03$ $-0.05$ $0.04$ $-0.07$ $0.04$
South $-0.05$ $0.03$ $-0.02$ $0.04$ $-0.07$ $0.04$
West $0.19$ $0.03$ $0.19$ $0.04$ $0.18$ $0.04$
US Born $-0.21$ $0.04$ $-0.33$ $0.04$ $-0.22$ $0.04$
Black $-0.28$ $0.04$ $-0.14$ $0.06$ $-0.18$ $0.05$
Other Races $0.05$ $0.03$ $0.01$ $0.04$ $-0.03$ $0.04$
Democrat $0.59$ $0.03$ $0.65$ $0.04$ $0.75$ $0.04$
Ind & Others $0.40$ $0.03$ $0.36$ $0.03$ $0.43$ $0.03$
Income/10000 (USD) $0.01$ $0.01$ $0.02$ $0.01$ $0.03$ $0.01$
$\delta_{1}$ $0.01$ $0.02$ $-0.10$ $0.02$ $-0.11$ $0.02$
$\delta_{2}$ $0.12$ $0.01$ $-0.00$ $0.02$ $0.04$ $0.02$
$\ln$ ML $-12415$ $-8301$ $-7914$
Confidence in charging infrastructure, assessed in the 2023 survey, is one of the strongest predictors of EV interest. Respondents who are very or extremely confident that the US will build the necessary charging infrastructure are 19.1 percentage points more likely to express strong interest in purchasing EVs and 18.8 percentage points less likely to report no interest compared to those who report lower confidence. This finding aligns with prior research, such as Carley et al. (2013) and Li et al. (2017). As shown in Table 1, 82% of respondents lack confidence that US will develop the infrastructure needed to support EV adoption. Taken together, these results suggest that skepticism about charging infrastructure is likely to be a significant barrier to broader adoption, highlighting the importance of both improving infrastructure and clearly communicating its availability, reliability, and expansion plans to the public.
Table 3: Covariate effects (CE) from ordinal probit models. CE for variables not statistically different from zero are marked with $\times$.
April 2021 May 2022 May–June 2023
$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
EV Info $-0.0851$ $-0.0388$ $0.0217$ $0.1022$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $-0.1955$ $-0.0565$ $0.0970$ $0.1551$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $-0.1880$ $-0.0553$ $0.0526$ $0.1908$
EV Owner $-0.1321$ $-0.0849$ $0.0051$ $0.2119$ $-0.1749$ $-0.0794$ $0.0231$ $0.2312$ $-0.1814$ $-0.0513$ $0.0421$ $0.1905$
GACC (TM) $0.1971$ $0.0453$ $-0.0945$ $-0.1478$ $0.3341$ $0.0133$ $-0.1507$ $-0.1968$ $0.3170$ $-0.0153$ $-0.1395$ $-0.1622$
GACC (RA) $0.1180$ $0.0391$ $-0.0490$ $-0.1082$ $0.1649$ $0.0267$ $-0.0639$ $-0.1278$ $0.1282$ $0.0079$ $-0.0486$ $-0.0875$
Age $<50$ $-0.0326$ $-0.0130$ $0.0099$ $0.0356$ $-0.0410$ $-0.0090$ $0.0129$ $0.0371$ $-0.0371$ $-0.0036$ $0.0126$ $0.0282$
Female $\times$ $\times$ $\times$ $\times$ $0.0268$ $0.0055$ $-0.0082$ $-0.0240$ $0.0301$ $0.0027$ $-0.0100$ $-0.0228$
Grad & Above $-0.0815$ $-0.0322$ $0.0276$ $0.0860$ $-0.1139$ $-0.0240$ $0.0378$ $0.1001$ $-0.1264$ $-0.0130$ $0.0445$ $0.0949$
Some College $-0.0373$ $-0.0152$ $0.0099$ $0.0426$ $-0.0541$ $-0.0118$ $0.0156$ $0.0503$ $-0.0517$ $-0.0052$ $-0.0166$ $0.0403$
Married ..absent... . ..absent... . ..absent... . ..absent... . $-0.0154$ $-0.0029$ $0.0050$ $0.0134$$\times$ $\times$ $\times$ $\times$
Metropolitan $-0.0534$ $-0.0179$ $0.0191$ $0.0522$ $-0.0662$ $-0.0099$ $0.0232$ $0.0528$ $-0.0512$ $-0.0027$ $0.0183$ $0.0356$
Midwest $0.0162$ $0.0061$ $-0.0051$ $-0.0172$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
South $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
West $-0.0378$ $-0.0164$ $0.0101$ $0.0441$ $-0.0457$ $-0.0109$ $0.0135$ $0.0431$ $-0.0454$ $-0.0053$ $0.0148$ $0.0360$
US Born $0.0416$ $0.0186$ $-0.0106$ $-0.0497$ $0.0777$ $0.0217$ $-0.0214$ $-0.0781$ $0.0551$ $0.0071$ $-0.0178$ $-0.0444$
Black $0.0614$ $0.0200$ $-0.0220$ $-0.0594$ $0.0349$ $0.0058$ $-0.0116$ $-0.0291$ $0.0460$ $0.0026$ $-0.0160$ $-0.0326$
Other Races $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Democrat $-0.1211$ $-0.0578$ $0.0403$ $0.1386$ $-0.1569$ $-0.0497$ $0.0513$ $0.1553$ $-0.1882$ $-0.0353$ $0.0677$ $0.1558$
Ind & Others $-0.0808$ $-0.0310$ $0.0219$ $0.0898$ $-0.0890$ $-0.0187$ $0.0266$ $0.0811$ $-0.1091$ $-0.0093$ $0.0356$ $0.0827$
$\Delta$I=\$$=\$50,000$$\times$ $\times$ $\times$ $\times$ $-0.0223$ $-0.0051$ $0.0066$ $0.0209$ $-0.0363$ $-0.0045$ $0.0116$ $0.0291$
Across all three surveys, previous EV ownership is consistently linked with a higher interest in purchasing an EV. Respondents who have owned an EV or hybrid vehicle are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest compared to those without prior ownership. Other studies such as Carley et al. (2013) and Buhmann and Criado (2023) have also documented this pattern, suggesting that firsthand experience reduces uncertainty and fosters positive perceptions of EVs. However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in the technology, making them more engaged with the EV market.
Respondents who think the government is doing too much to address climate change (GACC, TM) are 14.8–19.7 percentage points less likely to express strong interest in purchasing EVs and 19.7–33.4 percentage points more likely to report no interest. Those who believe the government is doing the right amount (GACC, RA) also express lower interest, though the effects are smaller.
Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs compared to older respondents. Our finding aligns with Zhao et al. (2022), Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025). For instance, Zhao et al. (2022) finds that EV buyers in Shanghai are predominantly younger, with most between the ages of 31 and 40.
Results from the 2022 and 2023 surveys also show that female respondents are modestly less likely to express strong interest and more likely to report no interest than males. The covariate effects on strong interest decrease between 2.3–2.4 percentage points; while the covariate effect for no interest increase between 2.7–3.0 percentage points. This finding is consistent several studies, including Carley et al. (2013), Krause et al. (2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men in their samples exhibit a stronger preference for EVs.
Higher education is also linked to greater EV interest. Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest compared to those with a high school diploma or less. Similar, though smaller, effects are observed for respondents with some college education. The positive relationship between education and EV interest is similarly documented in Carley et al. (2013), Krause et al. (2013), Buhmann and Criado (2023), and Mamkhezri (2025).
Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest. This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report EV preference to be higher among urban residents. Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest compared to those in the Northeast, though those in the Midwest and South show no significant difference from their Northeastern counterparts. This pattern likely reflects California’s leadership in EV adoption, driven in part by policies such as ZEV mandates and the Clean Vehicle Rebate Program, along with a robust charging network. [^5] See https://www.eia.gov/todayinenergy/detail.php?id=61082
US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest compared to non-US-born respondents, possibly reflecting differences in cultural attitudes, driving patterns, or perceptions about the cost of vehicle ownership. Such a difference in perception about EVs is also evident in the racial background of respondents. Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents, likely reflecting structural barriers, such as limited access to charging infrastructure (Hsu and Fingerman, 2021). Other non-white respondents show similar preferences to White respondents.
Political affiliation is a consistent and significant predictor of EV purchase intent. Democrats are 13.9–18.8 percentage points more likely to report strong interest in EVs and 12.1–18.8 percentage points less likely to report no interest compared to Republicans. Similarly, Independents and others are also about 8 percentage points more likely to express strong interest in purchasing EVs relative to Republicans. These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible compared to Republicans, as documented by Sintov et al. (2020). They also reflect the starkly contrasting policy approaches of the two major political parties, particularly during the tenures of Presidents Biden and Trump. During his first term (2017–2021), President Trump’s administration rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs. In contrast, the Biden administration adopted an aggressive pro-EV stance as part of its broader climate agenda, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
Conditional on other factors, an additional \$50,000 in income is either not significantly associated with EV interest or changes the probability of strong interest (no interest) by a meagre 2.1–2.9 (2.2–3.6) percentage points. Our results find support in Zhao et al. (2022) and Mamkhezri (2025), who report greater support for EVs among higher-income individuals. The small covariate effect for income may reflect the growing availability of EVs across various market segments and the presence of subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.
Overall, the results show that interest in EVs varies based on environmental attitudes, experience with EVs, confidence in charging infrastructure, and demographic factors. Some patterns likely reflect personal preferences, while others may reflect barriers such as limited awareness or charging access.
4.2 Results from Ordinal Quantile Models
We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to the probabilities of outcomes at specific quantiles, conditional on the covariates. In particular, Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models along with the logarithm of marginal likelihood, estimated separately for each survey data. Table 5 reports the corresponding covariate effects. The covariate effects at the 50th percentile align closely with those from the ordinal probit model. However, across all surveys, the covariate effects at the 20th and 80th quantiles are noticeably different compared to those from the ordinal probit model. As such, we find compelling evidence that the impact of covariates on the intent to purchase EV is not uniform, but varies across the distribution of EV purchase intent. This highlights the value of quantile modeling for capturing these variations. Below, we emphasize the key differences across quantiles.
While having heard or read extensively about EVs corresponds to higher purchase intent, the relationship is stronger among respondents more inclined toward adoption. In the 2021 survey and at the 80th percentile of EV purchase intent, respondents who have heard or read extensively about EVs are 12.6 percentage points more likely to express strong intent to purchase compared to those who have read or heard little to nothing. At the 50th and 20th percentiles of EV purchase intent, the difference is 9.9 and 6.1 percentage points, respectively. The smaller covariate effect of prior exposure to EV information at the 20th percentile, suggests that simply raising general awareness may not be enough to persuade consumers with low purchase interest. In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may be effective in influencing purchase decisions.
Similarly, the relationship between confidence that the US will build the necessary charging infrastructure (measured in the 2023 survey) and EV purchase intent varies across the distribution of EV purchase intent. At the 80th percentile, respondents who are confident in this development are 20.0 percentage points more likely to express strong intent to purchase compared to those who are not confident. The corresponding differences are 18.1 percentage points at the median and 13.3 percentage points at the 20th percentile. Conversely, at the 80th percentile, those confident in infrastructure development (relative to those who are not) are 14.5 percentage points less likely to express no interest, with differences of 18.3 and 18.6 percentage points at the median and 20th percentile, respectively. These findings suggest that confidence in infrastructure development strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent. Enhancing the affordability, accessibility, and reliability of charging infrastructure—and effectively communicating these efforts—can build consumer confidence and promote broader EV adoption.
Table 4: Results from quantile ordinal model: posterior mean (Mean) and standard deviation (Std) of the parameters, along with the logarithm of marginal likelihood, from April 2021, May 2022, and May–Jun 2023 survey data.
April 2021 May 2022 May–June 2023
20th 50th 80th 20th 50th 80th 20th 50th 80th
Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std
Intercept $-3.60$ $0.25$ $-0.58$ $0.17$ $2.60$ $0.25$ $-2.15$ $0.30$ $0.80$ $0.21$ $4.64$ $0.29$ $-2.81$ $0.30$ $-0.05$ $0.20$ $2.77$ $0.28$
EV Info $1.28$ $0.08$ $0.98$ $0.06$ $1.64$ $0.08$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $2.52$ $0.11$ $1.80$ $0.07$ $2.85$ $0.10$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $2.83$ $0.13$ $1.92$ $0.10$ $2.71$ $0.14$
EV Owner $2.80$ $0.13$ $1.76$ $0.10$ $2.91$ $0.16$ $2.94$ $0.16$ $1.83$ $0.11$ $3.31$ $0.20$ $2.65$ $0.16$ $1.99$ $0.11$ $3.08$ $0.19$
GACC (TM) $-2.81$ $0.13$ $-1.96$ $0.09$ $-2.49$ $0.13$ $-3.79$ $0.16$ $-2.63$ $0.10$ $-3.70$ $0.14$ $-4.00$ $0.17$ $-2.41$ $0.10$ $-3.45$ $0.14$
GACC (RA) $-1.79$ $0.10$ $-1.24$ $0.07$ $-1.55$ $0.11$ $-1.94$ $0.12$ $-1.48$ $0.08$ $-2.03$ $0.12$ $-1.77$ $0.12$ $-1.15$ $0.08$ $-1.48$ $0.12$
Age$<50$ $0.40$ $0.08$ $0.36$ $0.05$ $0.58$ $0.08$ $0.50$ $0.09$ $0.37$ $0.07$ $0.63$ $0.10$ $0.42$ $0.11$ $0.34$ $0.07$ $0.55$ $0.10$
Female $-0.11$ $0.08$ $-0.02$ $0.05$ $-0.03$ $0.07$ $-0.36$ $0.10$ $-0.22$ $0.06$ $-0.33$ $0.09$ $-0.49$ $0.10$ $-0.31$ $0.07$ $-0.27$ $0.09$
Grad & Above $1.42$ $0.14$ $-0.93$ $0.08$ $1.17$ $0.13$ $1.93$ $0.16$ $1.04$ $0.10$ $1.19$ $0.13$ $1.94$ $0.16$ $1.09$ $0.10$ $1.43$ $0.13$
Some College $0.66$ $0.14$ $0.47$ $0.09$ $0.57$ $0.12$ $1.08$ $0.16$ $0.53$ $0.10$ $0.49$ $0.13$ $0.92$ $0.16$ $0.42$ $0.10$ $0.53$ $0.13$
Married ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $0.22$ $0.13$ $0.12$ $0.08$ $0.19$ $0.12$ $0.02$ $0.13$ $0.08$ $0.08$ $0.12$ $0.11$
Metropolitan $0.77$ $0.12$ $0.64$ $0.08$ $0.71$ $0.12$ $0.91$ $0.16$ $0.59$ $0.10$ $0.79$ $0.13$ $0.79$ $0.16$ $0.46$ $0.10$ $0.61$ $0.13$
Midwest $-0.31$ $0.12$ $-0.19$ $0.08$ $-0.24$ $0.11$ $-0.09$ $0.15$ $-0.09$ $0.10$ $-0.11$ $0.14$ $-0.18$ $0.16$ $-0.15$ $0.10$ $-0.21$ $0.14$
South $-0.21$ $0.10$ $-0.12$ $0.07$ $-0.17$ $0.10$ $0.00$ $0.13$ $-0.02$ $0.09$ $-0.06$ $0.13$ $-0.22$ $0.15$ $-0.22$ $0.09$ $-0.17$ $0.13$
West $0.57$ $0.12$ $0.47$ $0.08$ $0.70$ $0.12$ $0.46$ $0.15$ $0.45$ $0.10$ $0.89$ $0.14$ $0.63$ $0.16$ $0.46$ $0.10$ $0.66$ $0.15$
US Born $-0.63$ $0.12$ $-0.49$ $0.09$ $-0.80$ $0.14$ $-0.89$ $0.15$ $-0.67$ $0.10$ $-1.25$ $0.16$ $-0.91$ $0.16$ $-0.45$ $0.11$ $-0.67$ $0.15$
Black $-1.23$ $0.15$ $-0.70$ $0.10$ $-0.56$ $0.14$ $-0.86$ $0.20$ $-0.25$ $0.14$ $-0.06$ $0.18$ $-0.94$ $0.16$ $-0.42$ $0.11$ $-0.36$ $0.16$
Other Races $0.08$ $0.10$ $0.14$ $0.07$ $0.25$ $0.11$ $-0.19$ $0.14$ $0.07$ $0.10$ $0.22$ $0.13$ $-0.30$ $0.15$ $0.00$ $0.10$ $0.02$ $0.14$
Democrat $2.00$ $0.11$ $1.33$ $0.08$ $1.94$ $0.12$ $2.30$ $0.14$ $1.45$ $0.09$ $1.90$ $0.13$ $2.49$ $0.14$ $1.71$ $0.10$ $2.40$ $0.14$
Ind & Others $1.29$ $0.10$ $0.88$ $0.07$ $1.42$ $0.11$ $1.32$ $0.12$ $0.80$ $0.08$ $1.06$ $0.11$ $1.40$ $0.13$ $0.95$ $0.08$ $1.44$ $0.12$
Income/10000 (USD) $0.04$ $0.01$ $0.02$ $0.01$ $0.00$ $0.01$ $0.07$ $0.02$ $0.05$ $0.01$ $0.05$ $0.02$ $0.12$ $0.02$ $0.06$ $0.01$ $0.08$ $0.02$
$\delta_{1}$ $1.16$ $0.02$ $0.91$ $0.02$ $1.45$ $0.02$ $1.09$ $0.02$ $0.76$ $0.02$ $1.25$ $0.02$ $1.14$ $0.02$ $0.74$ $0.02$ $1.21$ $0.02$
$\delta_{2}$ $1.53$ $0.02$ $0.95$ $0.02$ $1.35$ $0.02$ $1.46$ $0.02$ $0.84$ $0.02$ $1.16$ $0.02$ $1.52$ $0.02$ $0.93$ $0.02$ $1.19$ $0.02$
$\ln$ ML $-12871$ $-12529$ $-12901$ $-8563$ $-8357$ $-8624$ $-8237$ $-7981$ $-8194$
Table 5: Covariate effects (CE) from quantile analysis. CE for variables not statistically different from zero are marked with $\times$.
April 2021 Survey
20th Quantile 50th Quantile 80th Quantile
$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
EV Info $-0.0722$ $-0.0421$ $0.0527$ $0.0615$ $-0.0792$ $-0.0374$ $0.0172$ $0.0995$ $-0.0722$ $-0.0453$ $-0.0089$ $0.1265$
Env Better $-0.1958$ $-0.0191$ $0.1135$ $0.1014$ $-0.1771$ $-0.0730$ $0.1122$ $0.1379$ $-0.1384$ $-0.1127$ $0.0789$ $0.1722$
EV Owner $-0.1357$ $-0.0896$ $0.0633$ $0.1620$ $-0.1212$ $-0.0757$ $-0.0113$ $0.2083$ $-0.1053$ $-0.0953$ $-0.0265$ $0.2270$
GACC (TM) $0.2192$ $0.0097$ $-0.1217$ $-0.1072$ $0.1974$ $0.0635$ $-0.1203$ $-0.1406$ $0.1283$ $0.0779$ $-0.0575$ $-0.1487$
GACC (RA) $0.1216$ $0.0312$ $-0.0780$ $-0.0747$ $0.1097$ $0.0534$ $-0.0613$ $-0.1018$ $0.0763$ $0.0421$ $-0.0124$ $-0.1059$
Age $<50$ $-0.0237$ $-0.0106$ $0.0156$ $0.0187$ $-0.0303$ $-0.0134$ $0.0095$ $0.0343$ $-0.0265$ $-0.0164$ $0.0003$ $0.0427$
Metropolitan $-0.0483$ $-0.0174$ $0.0318$ $0.0338$ $-0.0565$ $-0.0225$ $0.0237$ $0.0553$ $-0.0345$ $-0.0185$ $0.0030$ $0.0500$
Midwest $0.0185$ $0.0080$ $-0.0123$ $-0.0141$ $0.0161$ $0.0068$ $-0.0053$ $-0.0177$ $0.0114$ $0.0066$ $-0.0003$ $-0.0177$
South $0.0124$ $0.0056$ $-0.0083$ $-0.0098$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
West $-0.0327$ $-0.0168$ $0.0220$ $0.0275$ $-0.0382$ $-0.0176$ $0.0092$ $0.0467$ $-0.0311$ $-0.0200$ $-0.0015$ $0.0526$
Black $0.0784$ $0.0256$ $-0.0519$ $-0.0520$ $0.0613$ $0.0244$ $-0.0259$ $-0.0598$ $0.0269$ $0.0141$ $-0.0014$ $-0.0396$
Democrat $-0.1242$ $-0.0523$ $0.0808$ $0.0958$ $-0.1081$ $-0.0631$ $0.0431$ $0.1280$ $-0.0828$ $-0.0676$ $0.0070$ $0.1434$
May 2022 Survey
EV Owner $-0.1782$ $-0.0723$ $0.0880$ $0.1624$ $-0.1573$ $-0.0695$ $0.0121$ $0.2147$ $-0.1465$ $-0.1110$ $-0.0071$ $0.2646$
GACC (TM) $0.3425$ $-0.0568$ $-0.1515$ $-0.1341$ $0.3304$ $0.0226$ $-0.1778$ $-0.1752$ $0.2471$ $0.0963$ $-0.1459$ $-0.1975$
GACC (RA) $0.1541$ $0.0026$ $-0.0790$ $-0.0777$ $0.1527$ $0.0497$ $-0.0852$ $-0.1171$ $0.1269$ $0.0392$ $-0.0308$ $-0.1352$
Age $<50$ $-0.0362$ $-0.0066$ $0.0199$ $0.0229$ $-0.0390$ $-0.0085$ $0.0130$ $0.0344$ $-0.0358$ $-0.0167$ $0.0072$ $0.0453$
Metropolitan $-0.0684$ $-0.0054$ $0.0360$ $0.0378$ $-0.0643$ $-0.0109$ $0.0258$ $0.0493$ $-0.0483$ $-0.0165$ $0.0114$ $0.0535$
Black $0.0642$ $0.0060$ $-0.0342$ $-0.0359$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Democrat $-0.1719$ $-0.0295$ $0.0927$ $0.1087$ $-0.1437$ $-0.0574$ $0.0601$ $0.1410$ $-0.0996$ $-0.0659$ $0.0217$ $0.1438$
May–June 2023 Survey
EV Infra $-0.1863$ $-0.0428$ $0.0956$ $0.1335$ $-0.1827$ $-0.0603$ $0.0616$ $0.1813$ $-0.1452$ $-0.0730$ $0.0187$ $0.1995$
EV Owner $-0.1684$ $-0.0350$ $0.0727$ $0.1307$ $-0.1842$ $-0.0542$ $0.0459$ $0.1925$ $-0.1580$ $-0.0817$ $0.0183$ $0.2215$
GACC (TM) $0.3323$ $-0.0635$ $-0.1405$ $-0.1283$ $0.3103$ $-0.0250$ $-0.1426$ $-0.1427$ $0.2512$ $0.0486$ $-0.1423$ $-0.1575$
GACC (RA) $0.1318$ $-0.0065$ $-0.0592$ $-0.0661$ $0.1238$ $0.0132$ $-0.0570$ $-0.0799$ $0.0975$ $0.0183$ $-0.0329$ $-0.0830$
Age $<50$ $-0.0296$ $-0.0007$ $0.0131$ $0.0173$ $-0.0363$ $-0.0034$ $0.0139$ $0.0258$ $-0.0349$ $-0.0082$ $0.0102$ $0.0329$
Metropolitan $-0.0568$ $0.0013$ $0.0250$ $0.0304$ $-0.0511$ $-0.0026$ $0.0204$ $0.0333$ $-0.0408$ $-0.0067$ $0.0129$ $0.0345$
Black $0.0665$ $-0.0010$ $-0.0296$ $-0.0359$ $0.0462$ $0.0029$ $-0.0182$ $-0.0309$ $0.0239$ $0.0039$ $-0.0069$ $-0.0209$
Democrat $-0.1832$ $-0.0084$ $0.0851$ $0.1066$ $-0.1815$ $-0.0384$ $0.0798$ $0.1401$ $-0.1393$ $-0.0660$ $0.0491$ $0.1562$
Beliefs about EVs’ environmental benefits (measured in the 2021 survey) also show varying relationships with purchase intent across quantiles. Among respondents at the 80th percentile, those who believe EVs are better for the environment are 17.2 percentage points more likely to express strong purchase intent compared to those at the same percentile who do not share this belief. The corresponding differences are 13.8 percentage points at the median and 10.1 percentage points at the 20th percentile. Conversely, such respondents are 13.8 percentage points less likely to express no interest at the 80th percentile, compared to 17.7 percentage points less likely at the median and 19.6 percentage points less likely at the 20th percentile. This finding suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters while easing skepticism among consumers with lower purchase interest; this indicates that highlighting benefits such as local air quality improvements could broaden support for EV adoption.
The influence of prior EV ownership also varies across quantiles. At the 80th percentile, respondents with prior EV ownership are 22.7 percentage points more likely to express strong purchase intent compared to those without prior ownership. This difference decreases to 20.8 percentage points at the median and 16.2 percentage points at the 20th percentile. A similar pattern is observed in both the 2022 and 2023 surveys. In the 2022 survey, the probability of strong interest increases by 26.5, 21.5, and 16.2 percentage points at the 80th, 50th, and 20th quantiles, respectively. For the same quantiles, the covariate effects for strong interest in the 2023 survey are 22.1, 19.2, and 13.1 percentage points.
When it comes to views on government action, respondents who believe the government is doing too much or enough to reduce climate change are significantly less likely to express strong interest in purchasing an EV, with this negative effect becoming more pronounced at higher percentiles. In the 2021 survey, those at the 80th percentile of EV purchase intent who hold this belief are 14.9 percentage points less likely to report strong purchase intent compared to those at the same percentile who do not. The gap is 14.1 percentage points at the median and 10.7 percentage points at the 20th percentile. Likewise, at the 80th percentile, these respondents are 12.8 percentage points more likely to express no interest in EVs, with the gap widening to 19.7 percentage points at the median and 21.9 percentage points at the 20th percentile. More recent surveys indicate even stronger covariate effects, with those who believe the government is doing too much on climate issues being approximately 33 percentage points more likely to express no interest in EVs at the 20th percentile and 25 percentage points more likely at the 80th percentile. These patterns suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonates across ideological lines. Rather than positioning EVs primarily as a climate solution, emphasizing other benefits—such as long-term cost savings and air quality improvements—may be more effective in engaging a broader range of potential buyers. Moreover, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns, particularly for consumers who are wary of climate-focused policies.
Figure 3: The figure presents the covariate effects for some selected variables from the April 2021 and May-June 2023 surveys, spanning the 5th to 95th quantiles with 5th quantile increments. Legend: ΔPr(Not at all $\Delta\Pr(\textrm{Not at all likely})$:  ; ΔPr(Not too $\Delta\Pr(\textrm{Not too likely})$:   ; ΔPr(Somewhat $\Delta\Pr(\textrm{Somewhat likely})$:   ; ΔPr(Very $\Delta\Pr(\textrm{Very likely})$:  .
Geographic and demographic factors show smaller but consistent relationships with EV purchase intent across all surveys. Younger respondents (under 50), metropolitan residents, and those living in the West are more likely to express higher purchase intent across all considered quantiles, with the strongest covariate effects observed at the 80th percentile.
Finally, the relationship between political party affiliation and EV purchase intent varies considerably across quantiles. Results from the 2021 survey show that at the 80th percentile Democrats are 14.3 percentage points more likely than Republicans to express strong purchase intent, with the difference narrowing to 9.6 percentage points among respondents at the 20th percentile. Similarly, at the 80th percentile Democrats are 8.3 percentage points less likely than Republicans to express no interest in EVs, while the difference was 12.4 percentage points at the 20th percentile. These patterns remain consistent in more recent surveys. The results indicate that political identity is closely linked to EV purchase intent across the entire distribution: at higher quantiles, Democrats are significantly more likely to express strong purchase intent, while at lower quantiles, they are far less likely to reject EVs outright. Emphasizing benefits that resonate across political ideologies—such as long-term cost savings and energy independence—may help broaden EV appeal across the political spectrum.
To provide a clearer understanding of how covariate effects vary across quantiles, Figure 3 presents the covariate effects for some selected variables from a series of ordinal quantile regression models, covering the 5th to the 95th percentiles of EV purchase intent in 5-percentile increments. Each panel corresponds to a different covariate; with the solid gray, dashed gray, dashed black, and solid black lines representing the change in probability of reporting “not likely at all,” “not too likely,” “somewhat likely,” and “very likely,” respectively, for a given change in covariate across the quantiles. The figure clearly illustrates the substantial differences in covariate effects across the distribution of EV purchase intent.
For example, the solid black line in these panels indicates that respondents at higher quantiles who have read or heard extensively about EVs, perceive them as environmentally beneficial, have previously owned an EV, identify as Democrats, or are confident in the United States’ ability to develop the necessary infrastructure are much more likely to report strong interest in purchasing EVs compared to those at lower quantiles. Conversely, the solid gray lines show that those who believe government climate efforts are excessive or adequate are significantly more likely to report no interest in buying EVs at lower quantiles and markedly less likely to report strong interest in buying EVs at higher quantiles.
Overall, these results highlight the need for tailored strategies to address the barriers to EV adoption for different population segments.
5 Conclusion
EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts. However, government support has varied between administrations. The Obama administration facilitated EV adoption by introducing federal tax credits for EV purchases and strengthening fuel economy and emissions standards, which incentivized automakers to develop electric models. In contrast, the Trump administration’s first term prioritized deregulation and rolled back these standards. The Biden administration implemented stricter emissions regulations and allocated substantial funding for EV infrastructure through the Bipartisan Infrastructure Law. A second Trump administration has signaled a likely return to deregulation, though specific policies remain unclear.
While policies and regulatory standards continue to evolve with changing administrations, the future of electric vehicles depends largely on the interest of consumers, which plays a crucial role in shaping national policies. Public opinion has a well-documented influence on US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has systematically analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of the covariates. In addition, no studies have applied quantile analysis to examine how different factors influence the distribution of intention to purchase EVs.
This study addresses these gaps by examining public attitudes toward EVs using three nationally representative surveys conducted by the Pew Research Center from 2021-2023. Given the ordinal nature of the response data, we employ both ordinal probit regression and ordinal quantile regression to model EV purchase intent as a function of four novel variables not considered in existing studies, prior EV ownership, political affiliation, and demographic characteristics such as age, gender, income, and geographic location. We adopt the Bayesian approach to modeling and estimation, as it offers the entire posterior distribution of the parameters, enables direct probability statements, and allows us to account for uncertainty in observations and parameters when evaluating the effects of covariates.
The results from ordinal probit models indicate that respondents who say they have read or heard a lot about EVs, perceive them as environmentally beneficial, have previously owned an EV or hybrid vehicle, and are confident in the US developing charging infrastructure are more likely to express strong interest in purchasing an EV. Other factors positively influencing the intention to purchase an EV include being under 50 years old, having a higher level of education, living in metropolitan areas, residing in the West (compared to the Northeast), identifying with the Democratic Party (rather than the Republican Party), and having a higher income. In contrast, those who believe government climate efforts are excessive or adequate, are female, live in the Midwest, were born in the US, or are Black (as opposed to White) are less likely to express strong interest and more likely to report no interest in EVs. Results from covariate effects reveal that prior EV ownership, confidence in EV infrastructure, belief in EVs’ environmental benefits, attitudes toward government climate policy, and political affiliation have the strongest influence on individuals’ decision to purchase EVs.
In addition, results from ordinal quantile analysis show that the relationship between these factors and EV purchase intent varies across the conditional distribution of the intention to purchase EVs. Respondents who are knowledgeable about EVs, believe in their environmental benefits, are confident in US charging infrastructure, have prior EV ownership, or identify as Democrats are especially likely to express strong purchase intent at higher quantiles of EV purchase intent. Similarly, those who believe EVs are environmentally beneficial, have confidence in infrastructure development, or identify as Democrats are particularly less likely to express no interest at lower quantiles. In contrast, respondents who believe the government is doing too much to address climate change are particularly more likely to express no interest at lower quantiles and particularly less likely to express strong interest at higher quantiles. Overall, the quantile models highlight significant variation in covariate effects across the distribution of EV purchase intent; thereby providing a more nuanced understanding than the ordinal probit model.
The findings demonstrate how targeted strategies may be able to help accelerate EV adoption. For consumers with low purchase intent (i.e., at lower quantiles), general awareness campaigns may have limited impact; instead, emphasizing practical aspects such as long-term cost savings, availability and reliability of charging infrastructure, and financial incentives could be more effective. For more interested consumers (i.e., at higher quantiles), reinforcing beliefs about environmental benefits can transform a likely customer to an actual EV owner. Confidence in charging infrastructure plays a dual role, both strengthening commitment among likely adopters and reducing outright rejection among skeptics, highlighting the potential value of visible investments and clear public communication about infrastructure progress. The results also indicate that political identity and views on climate policy are closely related to EV purchase intent, suggesting that framing EVs in terms of widely appealing features such as financial savings and energy independence—rather than solely as a climate solution—could resonate with a broader range of potential buyers. Moreover, for consumers wary of government climate policy, information provided by automakers, dealerships, and industry groups may carry more credibility than government-led campaigns. Overall, the empirical findings of our study provide useful insights for policymakers, industry stakeholders, and advertisers seeking to bolster adoption.